OpenStudy (anonymous):

Find the arc length of a central angle 300 degrees in a circle whose radius is 2 inches? A.10 π/ 3 in B.1200 in C. 22 π/ 3 in D. 75 in.

OpenStudy (bohotness):

which one do you think its the asnwer?

OpenStudy (anonymous):

Either a or c

OpenStudy (bohotness):

okay @ShadowLegendX

OpenStudy (bohotness):

i this one i need to check to make sure what i have is right before i can post it

OpenStudy (jhannybean):

\[\theta = \frac{s}{r}\]\[s= r\theta\]change \(\theta\) to radians. \[\theta =300^{\circ} \cdot \frac{\pi}{180^{\circ}} = \frac{5\pi}{3}\]\[s=\left(\frac{5\pi}{3}\right)(2)\]

OpenStudy (danjs):

^ @Jhannybean

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (triciaal):

use the ratio: angle / 360 = arc length/2*pi*r you are given the angle 300 and the radius, r = 2

OpenStudy (triciaal):

yes A